OpenStudy (anonymous):

Which answer is equal to the quotient in the expression below? (6x^6 - 6x^3 - 9x^2 + 3x) / (3x) A. 2x^3 - 2x^2 + 2x - 3 B. 2x^3 + 2x^2 + 2x C. 2x^5 - 2x^2 - 3x + 1 D. 2x^3 - 2x^2 - 3x - 2

OpenStudy (mokeira):

\[\frac{ 6x ^{6}-6x ^{3}-9x ^{2}+3x }{ 3x }\] factorize the numerator \[\frac{ 3x(2x ^{5}-2x ^{2}-3x+1) }{ 3x }\] the 3x cancel each other and you remain with \[2x ^{5}-2x ^{2}-3x+1\]

OpenStudy (mokeira):

that is choice C

OpenStudy (anonymous):

x^2-4-3(x-2)

OpenStudy (mokeira):

\[x ^{2}-4-3(x-2)=x ^{2}-4-3x+6\] \[x ^{2}+2-3x=(x-2)(x-1)\]

OpenStudy (mokeira):

@sieunhangao

OpenStudy (anonymous):

3x(x-7)+2x-14=0

OpenStudy (mokeira):

\[3x(x-7)+2x-14=0\] \[3x ^{2}-21x+2x-14=0\] \[3x ^{2}-19x-14=0\]

OpenStudy (anonymous):

lam sao nua

OpenStudy (mokeira):

meaning

OpenStudy (anonymous):

nghia la sao chi lun phan dap so di nha

OpenStudy (anonymous):

chi dum di

OpenStudy (mokeira):

English?

OpenStudy (anonymous):

tieng anh la sao

OpenStudy (anonymous):

nang ni pn chi minh lun phan tiep theo di

OpenStudy (anonymous):

hiu hu

OpenStudy (mokeira):

Tôi không nói được tiếng việt

OpenStudy (anonymous):

my next problem is to ask now

OpenStudy (anonymous):

3x(x-7)+2x-14=0

OpenStudy (mokeira):

\[3x(x-7)+2x-14=0\] First open up the brackets \[3x ^{2}-21x+2x-14=0\] Add the like terms \[3x ^{2}-20x-14=0\] What do you want to do?

OpenStudy (anonymous):

lam sao

OpenStudy (anonymous):

tim x

OpenStudy (anonymous):

nhanh len

OpenStudy (mokeira):

làm những gì bạn muốn làm với các phương trình? Bạn đã không được hướng dẫn

OpenStudy (triciaal):

@Mokeira the first time is correct the second time has an error -21 + 2 = 19 @sieunhangao the method is correct I hope you understand now.

OpenStudy (anonymous):

a^3b-ab^3+a^2+2ab+b^2

OpenStudy (anonymous):

giup toi bai do voi

OpenStudy (anonymous):

a^3b-ab^3+a^2+2ab+b^2